The ratio of magnetic field and magnetic moment at the centre of a current carrying circular loop is x. When both the current and
radius is doubled the ratio will be
(A) x / 8 (B) x / 4 (C) x / 2 (D) 2x

Grade:

2 Answers

Arun
3 years ago
The ratio of magnetic field and magnetic moment at the centre of a current carrying circular loop is x. When both the current and radius is doubled the ratio will be. x/8.
Vikas TU
3 years ago
Dear student 
B = mew *i/2a
B/M = x (given)
When both current and the radius re doubled 
mew /(2*pi*(2a)^3) = x/8